OpenStudy (anonymous):

If I have a problem like 3^-2 where the square root is negative, how do I solve it?

OpenStudy (anonymous):

Recall : a^-b=1/a^b

OpenStudy (kropot72):

3^-2 = 1/3^2 = 1/9

OpenStudy (anonymous):

How does it equal 1/3^2?

OpenStudy (anonymous):

I think it is just a definition... mathematicians decide that that's how it will be thought of, and it also happens to work with all the rules, so it is good...

OpenStudy (anonymous):

You have to wonder what does 3^-2 even mean, physically? 3^2 is easy to attach some meaning in geometry. But I think 3^-2 is slightly more strange. but 3^-2 = 1/3^2 works out well.

OpenStudy (kropot72):

Put 3^-2 over 1: 3^-2/1 Then multiply numerator and denominator by 3^2 remembering that (3^-2)*(3^2)=3^0 = 1. ((3^-2)*(3^2))/3^2 = 1/3^2

OpenStudy (anonymous):

Thanks to all that answered. Wish I could've given you all best answer :\

OpenStudy (kropot72):

You're welcome :)